# Rankers, Judges, and Assistants: Towards Understanding the Interplay of LLMs in Information Retrieval Evaluation

### v1

@@ 3.2.1 Vulnerability to Manipulation
A significant concern with the adoption of LLMs as judges is their potential vulnerability to adversarial manipulation. Initial research suggests that LLM judges might be vulnerable to keyword stuffing and other SEO strategies [2]. More broadly, knowledge of the (characteristics of the) LLM judge opens up ways to manipulate benchmarking results. This could lead to situations where a system achieves much higher scores under automatic evaluation with the LLM judge than under manual assessment [6]. This “eval hacking” undermines the purpose of evaluation, which is to accurately assess the true utility of a system for users.
@@
Understanding and Mitigating Vulnerabilities of LLM Judges While initial studies demonstrate the possibility of manipulating LLM judges, the extent of this vulnerability across different LLMs, attack strategies, and IR tasks remains largely unknown. What specific vulnerabilities do LLM judges exhibit, and how do these vulnerabilities vary across different models and evaluation settings? Furthermore, How can we design evaluation protocols that are robust to manipulation, ensuring that LLM-based evaluation remains a reliable and trustworthy measure of system performance? This is a crucial area for future research.
3.2.2 Systematic Biases
@@
A core challenge in using LLMs for both ranking and evaluation lies in the fundamental similarity of the two tasks: both involve estimating the relevance of a document to a given query. Several studies note the potential for significant systemic biases when LLMs are employed in both roles [43, 13, 54, 30]. In their summary of the LLM4IR workshop, Rahmani et al. [43] state “if we were to use an LLM both as an assessor and as a ranker, we could expect such a model to be favoured over other evaluated models.” Faggioli et al. [13] similarly caution that “if the model is used to judge relevance both for annotation and for retrieval, its evaluation would be overinflated, possibly with perfect performance.” If both ranking and automatic evaluation are predisposed towards certain types of results, it becomes difficult to identify truly relevant results. This can lead to the suppression of diverse perspectives and the promotion of homogenous content. Novel ranking approaches that deviate from the LLM’s inherent understanding of relevance might be unfairly penalized during the assessment phase. This phenomenon shares similarities with “reward hacking” observed in reinforcement learning, where agents exploit loopholes in the reward function to achieve high scores without genuinely solving the underlying task [4]. A particularly concerning form of this bias is circularity, where retrieval models are trained on LLM-generated labels [13, 43, 6]. This creates a self-reinforcing loop, where the ranker learns to produce outputs that the LLM judge deems relevant, further amplifying any existing biases.
@@
Interrelated Systemic Biases in LLM-Based Evaluation While the potential for systemic biases in LLM-based IR evaluation is acknowledged, the specific interactions and magnitudes of these biases remain largely unquantified. We identify three interrelated potential biases:
@@
  • Bias Towards LLM-Generated Text: LLM judges might exhibit an inherent preference for text generated by LLMs, regardless of the ranking system that retrieved it. This could be due to factors like stylistic similarities, reduced noise, or other characteristics of LLM-generated text. Indeed, studies have observed that LLMs exhibit bias favoring texts generated by the same underlying model [27, 37]. However, there is a significant lack of studies systematically quantifying the extent to which LLM judges favor LLM-generated text in the specific context of IR evaluation.
  • Combined Bias (LLM Ranker + LLM-Generated Text): The most complex scenario involves the potential interaction of the two biases above. Dai et al. [9] show that neural retrievers prefer LLM-generated content, but their analysis relies on human judgments, not LLM judges. Does an LLM judge exhibit an even stronger preference for LLM-generated text when it is retrieved by an LLM-based ranker? This synergistic effect, if present, could significantly distort evaluation outcomes.
@@ 4 Experiments
4.1 Experiment Design
We study the classic ad hoc retrieval task where a ranked list of documents are returned in response to a user query. We follow a standard retrieve-then-rerank paradigm, where an initial set of potentially relevant documents is identified by a fast and efficient first stage retriever, which are then subsequently reranked by a computationally more intensive but more accurate model. Our focus lies specifically in this reranking stage, noting that LLMs may also be used for retrieval [53].
@@ 4.1 Experiment Design
4.2 Experimental Setup
We utilize the TREC Deep Learning (DL) 2019 and 2020 datasets [7, 8], chosen due to their extensive use in prior research in this area. Both use the MS MARCO v1 passage corpus, which contains 8.8 million passages. We adopt the convention of referring to passages as “documents,” even if the unit of retrieval are passages in our experiments. The two datasets contain 43 and 54 queries, respectively, with human relevance annotations by TREC assessors.
Following [41, 51], all comparisons are based on the reranking of the top 100 passages retrieved by BM25 [22]. To ensure a fair comparison between human and LLM judges, we filter results that have not been judged by TREC assessors (instead of treating them as non-relevant). For simplicity, we report only on NDCG@10, which is the official evaluation metric of the DL track.
4.2.1 LLM Judges
For automatic assessment, we use two model generations of a powerful commercial LLM, Gemini, in two sizes within each generation: v1 Nano, v1 Pro, v1.5 Flash, and v1.5 Pro. We use the best applicable prompt [^1] We use the prompt that considers multiple aspects (A), but not role (R) nor multiple judges (M); narrative (N) and description (D) are unavailable for TREC DL. in [54] based on the open source implementation UMBRELA [56], with judgments performed on a 4-point scale. We set top-p =1 and the temperature to 0.
4.2.2 LLM Rankers
@@
  • RankT5 [64] is a reranker that uses T5 [42] and listwise ranking loss during supervised fine-tuning. It is considered a state-of-the-art supervised LLM-based ranker.
  • RG [19] is a pointwise prompting method based on Relevance Generation, where the prompt asks “Does the passage answer the query?” and the logit of “Yes” is used as the ranking score. We test RG with FLAN-T5-XXL and FLAN-UL2. Note that RG requires internal logits of output tokens and thus cannot be used with black-box LLMs such as Gemini.
  • PRP [41] is a pairwise prompting approach that is effective and robust across LLMs with different sizes. Given a query and two passages, the prompt asks “Which of the two passages is more relevant to the query?” The winning rate is used as the ranking score for each passage.
Figure 2: Illustration of oracle rankers, ordered by their expected performance, assuming that the top three results are highly relevant and the bottom three are non-relevant.
Table 1: Results (NDCG@10) on the TREC DL 2019 and 2020 collections using both human and LLM judges. The best LLM and Oracle reranking approaches per judge are boldfaced.
@@
Figure 3: Visualization of the performance of selected rankers from Table 1.
4.2.3 Oracle Rankers
@@
4.2.4 Measuring Alignment
Following prior work (cf. Section 3.1.1) we measure agreement with TREC judges (on all human-judged query-document pairs) in terms of Cohen’s $\kappa$ using both graded and binary relevance labels. Following Faggioli et al. [13], we create binary relevance labels by merging levels 0 and 1 (non-relevant) and levels 2 and 3 (relevant). Additionally, we report on relative system ordering in terms of Kendall’s $\tau$.
Table 2: Agreement between LLM and human judges (1) on individual relevance judgments (Cohen’s $\kappa$) using both graded and binary labels and (2) on relative ordering of systems (Kendall’s $\tau$) considering all systems in Table 1 and Oracle rankers only.
@@ 4.3 Results
Choice of LLM How well do LLM-based judgments align with human assessments when using different variations of LLM judges from the same model family? Looking at the evaluation scores of various rankers, we observe generally good agreement among the three largest models. In terms of agreement with human judges on individual relevance judgments (see Cohen’s $\kappa$ Table 2) the results are comparable to those reported in prior work for these datasets [56], with the newer v1.5 models performing clearly better than the v1 models. Interestingly, within this newer model generation, a larger model is not necessarily more capable, at least not according to this measure; v1.5 Flash shows much better agreement with humans when a graded relevance scale is used than the v1.5 Pro. On the other hand, the smallest LLM (v1 Nano) is unable to provide useful judgments, as evidenced by the Cohen’s $\kappa$ values being close to 0. While this model may be capable in other tasks [15], our results clearly indicate its unsuitability for judging relevance in this specific context. Therefore, we exclude the v1 Nano judge from subsequent analyses and discussions referring to “LLM judges.”
@@
Bias Towards LLM-generated Text Do LLM judges exhibit biases towards LLM-generated text (independent of the ranking mechanism used to retrieve that text)? We investigate this by comparing LLM judge assessments of original human-written documents and their AI-assisted counterparts. Using the MS MARCO dataset, which predates the widespread adoption of modern AI writing tools, we can reasonably assume that the original documents represent content created without significant AI assistance. To avoid bias potentially introduced by initial retrieval, we take a balanced sample: for each year (DL19 and DL20) we randomly sample 500 query-document pairs for each of the four relevance levels, resulting in a total of 4000 query-document pairs. We refer to this set as Original. We then employ our second most capable LLM (Gemini v1.5 Flash) to create an AI-rewritten version of each document in the Original set, following the methodology of Dai et al. [9]. [^2] They use the straightforward prompt “Please rewrite the following text: human-written text” in a zero-shot setting. This rewritten set is referred to as Rewritten. We shall assume that this rewriting process does not substantially alter the relevance of the documents to their corresponding queries, as verified by human assessors in [9]. Figure 4 presents the results using our most capable LLM (Gemini v1.5 Pro) as the judge. We can observe on the Original data that the LLM judge is lenient in its assessment of relevance, and specifically in labeling non-relevant documents as partially relevant. However, the judge does not appear to systematically inflate scores for the highest relevance level. Crucially, when comparing these results to the judgments on the Rewritten (LLM-generated) text, we do not observe a distributional shift towards higher relevance levels. In fact, the Rewritten documents show a slight increase in lower relevance labels. While these findings are specific to this particular combination of LLM rewriter and judge, they provide evidence against a general bias towards LLM-generated content, even when both models are from the same family.
Figure 4: Relevance levels estimated by an LLM judge (Gemini v1.5 Pro) for Original vs. AI-assisted content (Rewrites, using Gemini v1.5 Flash). According to human assessors, the labels should be uniformly distributed across the four relevance classes, as indicated by the dashed horizontal line.
@@
Figure 5: Distribution of score differences of a RankT5 ranker on LLM-rewritten vs. original text on a sample of 4000 query-document pairs.
5 The Role and Challenges of LLM Judges in IR
@@
Clarifying the Purpose Most studies, albeit often implicitly, employ LLMs as judges with the aim of replacing human assessors. However, as we discuss below, fundamental limitations preclude this possibility. We argue (in line with [30, 13]) that a more appropriate and productive goal should be to enable more effective use of limited human assessor time and resources. This shift in perspective—from replacement to reducing human effort—is crucial for guiding the development and deployment of LLM judges.
Acknowledging Fundamental Limitations It is essential to recognize that both ranking and relevance assessment address the same problem: predicting the relevance of a document to a given query. This inherent overlap introduces fundamental limitations when using LLMs for both tasks. Clarke and Dietz [6] argue that “A true gold standard must originate from human assessments, as only humans can determine the relevance of information in a way that reflects real-world utility.” Soboroff [48] highlights the inherent uncertainty of relevance, noting that it is “defined with respect to the entire cognitive state of the person, which changes as they use the system.” Because of these inherent limitations, “LLM assessments may themselves represent a strong ranking method, rather than a valid evaluation metric” [6]. Recent work, such as [36], has begun to provide uncertainty measures with LLM relevance predictions.
@@
5.1 Guidelines for Employing LLMs as Judges
@@
  • Consistent Evaluation Across Systems. All systems being compared within a single evaluation should be assessed using the same LLM judge configuration (model, prompt, settings). This ensures a fair and unbiased comparison, avoiding situations where some systems are evaluated with a more lenient or biased judge than others. Specifically, LLM judges should not be used selectively to fill “relevance holes” in existing human judgments [1].
@@
  • Employing Multiple LLMs as Judges. Using a combination of different LLM judges can help mitigate biases stemming from LLMs favoring responses from their own model family and improve robustness (see, e.g., [18]). Reporting the distribution of scores across different judges, as suggested by Rahmani et al. [43], can further enhance robustness.
  • Alignment with Human Preferences. Ensuring alignment between human and LLM raters is a substantial effort that needs to be continuously monitored and refined [54]. Ideally, results reported on LLM judges should also include human validation of the results on a representative sample. Researchers should also exercise care when making research claims based on results from LLM judges.
5.2 Open Questions and Future Directions

### v2

@@ 3.2.1. Vulnerability to Manipulation
A significant concern with the adoption of LLMs as judges is their potential vulnerability to adversarial manipulation. Initial research suggests that LLM judges might be vulnerable to keyword stuffing and other SEO strategies (Alaofi et al., 2024). More broadly, knowledge of the (characteristics of the) LLM judge opens up ways to manipulate benchmarking results. This could lead to situations where a system achieves much higher scores under automatic evaluation with the LLM judge than under manual assessment (Clarke and Dietz, 2024). This “eval hacking” undermines the purpose of evaluation, which is to accurately assess the true utility of a system for users.
@@
Issue #3: Understanding and Mitigating Vulnerabilities of LLM Judges While initial studies demonstrate the possibility of manipulating LLM judges, the extent of this vulnerability across different LLMs, attack strategies, and IR tasks remains largely unknown. What specific vulnerabilities do LLM judges exhibit, and how do these vulnerabilities vary across different models and evaluation settings? Furthermore, How can we design evaluation protocols that are robust to manipulation, ensuring that LLM-based evaluation remains a reliable and trustworthy measure of system performance? This is a crucial area for future research.
3.2.2. Systematic Biases
@@
A core challenge in using LLMs for both ranking and evaluation lies in the fundamental similarity of the two tasks: both involve estimating the relevance of a document to a given query. Several studies note the potential for significant systemic biases when LLMs are employed in both roles (Rahmani et al., 2024; Faggioli et al., 2023; Thomas et al., 2024; MacAvaney and Soldaini, 2023). In their summary of the LLM4IR workshop, Rahmani et al. (2024) state “if we were to use an LLM both as an assessor and as a ranker, we could expect such a model to be favoured over other evaluated models.” Faggioli et al. (2023) similarly caution that “if the model is used to judge relevance both for annotation and for retrieval, its evaluation would be overinflated, possibly with perfect performance.” If both ranking and automatic evaluation are predisposed towards certain types of results, it becomes difficult to identify truly relevant results. This can lead to the suppression of diverse perspectives and the promotion of homogenous content. Novel ranking approaches that deviate from the LLM’s inherent understanding of relevance might be unfairly penalized during the assessment phase. This phenomenon shares similarities with “reward hacking” observed in reinforcement learning, where agents exploit loopholes in the reward function to achieve high scores without genuinely solving the underlying task (Chen et al., 2024b). A particularly concerning form of this bias is circularity, where retrieval models are trained on LLM-generated labels (Faggioli et al., 2023; Rahmani et al., 2024; Clarke and Dietz, 2024). This creates a self-reinforcing loop, where the ranker learns to produce outputs that the LLM judge deems relevant, further amplifying any existing biases.
@@
Issue #4: Interrelated Systemic Biases in LLM-Based Evaluation While the potential for systemic biases in LLM-based IR evaluation is acknowledged, the specific interactions and magnitudes of these biases remain largely unquantified. We identify three interrelated potential biases:
@@
  • Bias Towards LLM-Generated Text: LLM judges might exhibit an inherent preference for text generated by LLMs, regardless of the ranking system that retrieved it. This could be due to factors like stylistic similarities, reduced noise, or other characteristics of LLM-generated text. Indeed, studies have observed that LLMs exhibit bias favoring texts generated by the same underlying model (Liu et al., 2024b; Panickssery et al., 2024). However, there is a significant lack of studies systematically quantifying the extent to which LLM judges favor LLM-generated text in the specific context of IR evaluation.
  • Combined Bias (LLM Ranker + LLM-Generated Text): The most complex scenario involves the potential interaction of the two biases above. Dai et al. (2024b) show that neural retrievers prefer LLM-generated content, but their analysis relies on human judgments, not LLM judges. Does an LLM judge exhibit an even stronger preference for LLM-generated text when it is retrieved by an LLM-based ranker? This synergistic effect, if present, could significantly distort evaluation outcomes.
@@ 4. Experiments
4.1. Experiment Design
We study the classic ad hoc retrieval task where a ranked list of documents are returned in response to a user query. We follow a standard retrieve-then-rerank paradigm, where an initial set of potentially relevant documents is identified by a fast and efficient first stage retriever, which are then subsequently reranked by a computationally more intensive but more accurate model. Our focus lies specifically in this reranking stage, noting that LLMs may also be used for retrieval (Tay et al., 2022).
@@ 4.1. Experiment Design
4.2. Experimental Setup
We utilize the TREC Deep Learning (DL) 2019 and 2020 datasets (Craswell et al., 2019, 2020), chosen due to their extensive use in prior research in this area. Both use the MS MARCO v1 passage corpus, which contains 8.8 million passages. We adopt the convention of referring to passages as “documents,” even if the unit of retrieval are passages in our experiments. The two datasets contain 43 and 54 queries, respectively, with human relevance annotations by TREC assessors.
Following (Qin et al., 2024; Sun et al., 2023), all comparisons are based on the reranking of the top 100 passages retrieved by BM25 (Lin et al., 2021). To ensure a fair comparison between human and LLM judges, we filter results that have not been judged by TREC assessors (instead of treating them as non-relevant). For simplicity, we report only on NDCG@10, which is the official evaluation metric of the DL track.
4.2.1. LLM Judges
For automatic assessment, we use two model generations of a powerful commercial LLM, Gemini, in two sizes within each generation: v1 Nano, v1 Pro, v1.5 Flash, and v1.5 Pro. We use the best applicable prompt [^1] We use the prompt that considers multiple aspects (A), but not role (R) nor multiple judges (M); narrative (N) and description (D) are unavailable for TREC DL. in (Thomas et al., 2024) based on the open source implementation UMBRELA (Upadhyay et al., 2024b), with judgments performed on a 4-point scale. We set top-p =1 and the temperature to 0.
4.2.2. LLM Rankers
@@
  • RankT5 (Zhuang et al., 2023) is a reranker that uses T5 (Raffel et al., 2020) and listwise ranking loss during supervised fine-tuning. It is considered a state-of-the-art supervised LLM-based ranker.
  • RG (Liang et al., 2023) is a pointwise prompting method based on Relevance Generation, where the prompt asks “Does the passage answer the query?” and the logit of “Yes” is used as the ranking score. We test RG with FLAN-T5-XXL and FLAN-UL2. Note that RG requires internal logits of output tokens and thus cannot be used with black-box LLMs such as Gemini.
  • PRP (Qin et al., 2024) is a pairwise prompting approach that is effective and robust across LLMs with different sizes. Given a query and two passages, the prompt asks “Which of the two passages is more relevant to the query?” The winning rate is used as the ranking score for each passage.
Figure 2. Illustration of oracle rankers, ordered by their expected performance, assuming that the top three results are highly relevant and the bottom three are non-relevant.
Table 1. Results (NDCG@10) on the TREC DL 2019 and 2020 collections using both human and LLM judges. The best LLM and Oracle reranking approaches per judge are boldfaced.
@@
Figure 3. Visualization of the performance of selected rankers from Table 1.
4.2.3. Oracle Rankers
@@
4.2.4. Measuring Alignment
Following prior work (cf. Section 3.1.1) we measure agreement with TREC judges (on all human-judged query-document pairs) in terms of Cohen’s $\kappa$ using both graded and binary relevance labels. Following Faggioli et al. (2023), we create binary relevance labels by merging levels 0 and 1 (non-relevant) and levels 2 and 3 (relevant). Additionally, we report on relative system ordering in terms of Kendall’s $\tau$.
Table 2. Agreement between LLM and human judges (1) on individual relevance judgments (Cohen’s $\kappa$) using both graded and binary labels and (2) on relative ordering of systems (Kendall’s $\tau$) considering all systems in Table 1 and Oracle rankers only.
@@ 4.3. Results
Choice of LLM How well do LLM-based judgments align with human assessments when using different variations of LLM judges from the same model family? Looking at the evaluation scores of various rankers, we observe generally good agreement among the three largest models. In terms of agreement with human judges on individual relevance judgments (see Cohen’s $\kappa$ Table 2) the results are comparable to those reported in prior work for these datasets (Upadhyay et al., 2024b), with the newer v1.5 models performing clearly better than the v1 models. Interestingly, within this newer model generation, a larger model is not necessarily more capable, at least not according to this measure; v1.5 Flash shows much better agreement with humans when a graded relevance scale is used than the v1.5 Pro. On the other hand, the smallest LLM (v1 Nano) is unable to provide useful judgments, as evidenced by the Cohen’s $\kappa$ values being close to 0. While this model may be capable in other tasks (Gemini Team Google, 2023), our results clearly indicate its unsuitability for judging relevance in this specific context. Therefore, we exclude the v1 Nano judge from subsequent analyses and discussions referring to “LLM judges.”
@@
Bias Towards LLM-generated Text Do LLM judges exhibit biases towards LLM-generated text (independent of the ranking mechanism used to retrieve that text)? We investigate this by comparing LLM judge assessments of original human-written documents and their AI-assisted counterparts. Using the MS MARCO dataset, which predates the widespread adoption of modern AI writing tools, we can reasonably assume that the original documents represent content created without significant AI assistance. To avoid bias potentially introduced by initial retrieval, we take a balanced sample: for each year (DL19 and DL20) we randomly sample 500 query-document pairs for each of the four relevance levels, resulting in a total of 4000 query-document pairs. We refer to this set as Original. We then employ our second most capable LLM (Gemini v1.5 Flash) to create an AI-rewritten version of each document in the Original set, following the methodology of Dai et al. (2024b). [^2] They use the straightforward prompt “Please rewrite the following text: human-written text” in a zero-shot setting. This rewritten set is referred to as Rewritten. We shall assume that this rewriting process does not substantially alter the relevance of the documents to their corresponding queries, as verified by human assessors in (Dai et al., 2024b). Figure 4 presents the results using our most capable LLM (Gemini v1.5 Pro) as the judge. We can observe on the Original data that the LLM judge is lenient in its assessment of relevance, and specifically in labeling non-relevant documents as partially relevant. However, the judge does not appear to systematically inflate scores for the highest relevance level. Crucially, when comparing these results to the judgments on the Rewritten (LLM-generated) text, we do not observe a distributional shift towards higher relevance levels. In fact, the Rewritten documents show a slight increase in lower relevance labels. While these findings are specific to this particular combination of LLM rewriter and judge, they provide evidence against a general bias towards LLM-generated content, even when both models are from the same family.
Figure 4. Relevance levels estimated by an LLM judge (Gemini v1.5 Pro) for Original vs. AI-assisted content (Rewrites, using Gemini v1.5 Flash). According to human assessors, the labels should be uniformly distributed across the four relevance classes, as indicated by the dashed horizontal line.
@@
Figure 5. Distribution of score differences of a RankT5 ranker on LLM-rewritten vs. original text on a sample of 4000 query-document pairs.
@@
5. The Role and Challenges of LLM Judges in IR
@@
Clarifying the Purpose Most studies, albeit often implicitly, employ LLMs as judges with the aim of replacing human assessors. However, as we discuss below, fundamental limitations preclude this possibility. We argue (in line with (MacAvaney and Soldaini, 2023; Faggioli et al., 2023)) that a more appropriate and productive goal should be to enable more effective use of limited human assessor time and resources. This shift in perspective—from replacement to reducing human effort—is crucial for guiding the development and deployment of LLM judges.
Acknowledging Fundamental Limitations It is essential to recognize that both ranking and relevance assessment address the same problem: predicting the relevance of a document to a given query. This inherent overlap introduces fundamental limitations when using LLMs for both tasks. Clarke and Dietz (2024) argue that “A true gold standard must originate from human assessments, as only humans can determine the relevance of information in a way that reflects real-world utility.” We must also recognize that relevance itself carries an intrinsic uncertainty; it depends on the entire cognitive state of the person, which changes as they use the system (Saracevic, 1996). Because of these inherent limitations, “LLM assessments may themselves represent a strong ranking method, rather than a valid evaluation metric” (Clarke and Dietz, 2024). Recent work, such as (Oosterhuis et al., 2024), has begun to provide uncertainty measures with LLM relevance predictions.
@@
5.1. Guidelines for Employing LLMs as Judges
@@
  • Consistent Evaluation Across Systems. All systems being compared within a single evaluation should be assessed using the same LLM judge configuration (model, prompt, settings). This ensures a fair and unbiased comparison, avoiding situations where some systems are evaluated with a more lenient or biased judge than others. Specifically, LLM judges should not be used selectively to fill “relevance holes” in existing human judgments (Abbasiantaeb et al., 2024).
@@
  • Employing Multiple LLMs as Judges. Using a combination of different LLM judges can help mitigate biases stemming from LLMs favoring responses from their own model family and improve robustness (see, e.g., (Jacovi et al., 2025)). Reporting the distribution of scores across different judges, as suggested by Rahmani et al. (2024), can further enhance robustness.
  • Alignment with Human Preferences. Ensuring alignment between human and LLM raters is a substantial effort that needs to be continuously monitored and refined (Thomas et al., 2024). Ideally, results reported on LLM judges should also include human validation of the results on a representative sample. Researchers should also exercise care when making research claims based on results from LLM judges.
5.2. Open Questions and Future Directions